The cost of producing 60 small fiberglass sailboats per year, and the cost of producing sails and fittings necessary to make the boats seaworthy in a single plant, are together R780,000. If produced in separate plants, the boats would cost R540,000, and the sails and fittings would cost R180,000. From this information, what can be learned about (1) economies of scale and (2) economies of scope in the production of sailboats, sails, and fittings? Perform any necessary calculations and explain. [3


1
Expert's answer
2022-02-22T12:08:13-0500

Solution:

1) When produced separately, the average cost of production is

540000/60 + 180000/60 = 12000

When produced jointly, the average cost of production is 780000/60 = 13000.

Since joint production increases the average cost, there are diseconomies of scale.

2) When produced separately, the total cost of production is 540000 + 180000 = 720000.

When produced jointly, the total cost of production is 780000.

Since joint production increases the total cost, there are diseconomies of scope